OpenStudy (princess13jb):

Please help! I don't know how to solve this question! A shuffle board disk is accelerated to a speed of 5.5 m/s and released. If the coefficient of kinetic friction between the disk and concrete is 0.34, how far does the disk go before it comes to a stop?

OpenStudy (harman.singh):

1) Solve for the acceleration \[a=\mu \times g\] 2) Use the constant acceleration formula to solve for the distance \[v^2=u^2+2ax\] - v is the final velocity - u is the initial velocity - a is the acceleration(found previously) - x is the distance

OpenStudy (princess13jb):

I haven't learned that formula yet

OpenStudy (harman.singh):

The first one?

OpenStudy (princess13jb):

Both

OpenStudy (harman.singh):

What sort of formulas have been using previously to this question?

OpenStudy (princess13jb):

I use Fg=mg Ff= uFn a= (Fa-Ff)/m F=ma for all my friction problems

OpenStudy (harman.singh):

The first formula I showed above actually comes from the formula 'Ff=uFn'. So you can use it in your question. \[F _{f}=\mu \times F _{n}\] - The frictional force(Ff) for an object moving linearly can be found by multiplying its mass with acceleration. - The Normal force(Fn) is equal to the Weight force which is the object's mass multiplied by gravity. Once we substitute this information into our formula, we get: \[ma = \mu \times mg\] Notice that m is common on both sides of the equation and we can cancel it out. So we are left with: \[a=\mu \times g\] and thats where that formula comes from.

OpenStudy (princess13jb):

O ok, but how do I find the distance?

OpenStudy (harman.singh):

For distance, use the second formula I showed you. \[v^2=u^2+2ax\] - v is the final velocity. It says in the question that the disk comes to a stop. So the final velocity(v) is equal to 0 - u is the initial velocity given to you in the question - a is the acceleration which comes from the formula a=ug - x is the distance that you are trying to find. Rearrange the equation to solve for x.

OpenStudy (princess13jb):

Thank you for helping

OpenStudy (harman.singh):

My pleasure!

OpenStudy (mousejkl):

This

OpenStudy (mousejkl):

I need to figure this out help me out

OpenStudy (mousejkl):

i give you a meadel